Rules and Properties of Probability

Complement Rule

For any event A, P(A)=1−P(A′)

Proof
By definition of A′, A∪A′=S so then since A,A′ are disjoint we have:

1=P(S)=P(A∪A′)=P(A)+P(A′)

☐

A Series System

Consider a series-connected system of 5 components. Say A is the event that the system fails. For A to happen, just one component must fail. We have:

A={FFFFF,FFFFS,...,FSSSS}=(A′)′={SSSSS}′

Clearly since P(A′)=125=132 then P(A)=3132.

P(A)≤1

For any event A, P(A)≤1.

Proof

Using the complement rule: 1=P(A)+P(A′)≥P(A) since P(A′)≥0 by Axiom 2.
☐

Addition Rule

For any events A,B:

P(A \cup B) = P(A) + P(B) - P(A \cap B) $${ #70bda0}

Proof
Note that A∪B=A∪(B∩A′), which is used without too much proof.
Pasted image 20240924092909.png

Because A and B∩A′ are disjoint, then P(A∪B)=P(A)+P(B∩A′) using the complement rule. But B=(B∩A)∪(B∩A′), so then since B∩A and B∩A′ are disjoint then P(B)=P(B∩A)+P(B∩A′). Combining these two:

P(A∪B)=P(A)+P(B∩A′)=P(A)+(P(B)−P(B∩A))

☐
These properties generally extend for 3 or more sets as well. This patter of additions and subtractions is called the inclusion-exclusion principle:

P(A∪B∪C)=P(A)+P(B)+P(C)−P(A∩B)−P(A∩C)−P(B∩C)+P(A∩B∩C)

Example

Residential Probabilities

If a suburb has 60% of residents with internet service, 80% with television service, and 50% with both, what is the probability that a randomly selected household has at least one of these services? Probability of exactly one service?

Proof
Here let I be the internet service event, and T for television. Here P(I)=0.6 and P(T)=0.8. Consequently P(I∩T)=0.5. Using the addition rule:

P(at least one service)=P(I∪T)=P(I)+P(T)−P(I∩T)=0.6+0.8−0.5=0.9

Similarly for exactly one service:

P(exactly one service)=P(I∩T′)+P(I′∩T)P(I∩T′)+P(I∩T)+P(I′∩T)+P(no service)=1P(exactly one)+P(I∩T)+P(I′∩T′)=1P(exactly one)+0.5+0.1=1⇒P(exactly one)=0.4

☐

References

  1. [[Matthew A. Carlton, Jay L. Devore - Probability with STEM Applications-Wiley (2020).pdf#page=46]]